The internal rate of return method is used by Testerman Construction Co. in analyzing a capital expenditure proposal that involves an investment of $41,424 and annual net cash flows of $16,000 for each of the four years of its useful life.

Present Value of an Annuity of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 1.833 1.736 1.690 1.626 1.528
3 2.673 2.487 2.402 2.283 2.106
4 3.465 3.170 3.037 2.855 2.589
5 4.212 3.791 3.605 3.352 2.991
6 4.917 4.355 4.111 3.784 3.326
7 5.582 4.868 4.564 4.160 3.605
8 6.210 5.335 4.968 4.487 3.837
9 6.802 5.759 5.328 4.772 4.031
10 7.360 6.145 5.650 5.019 4.192

a. Determine a present value factor for an annuity of $1 which can be used in determining the internal rate of return. If required, round your answer to three decimal places.

___________

b. Using the factor determined in part (a) and the present value of an annuity of $1 table above, determine the internal rate of return for the proposal.
________ %
Answer #1

Let irr be x%
At irr,present value of inflows=present value of outflows.

41,424=16000/1.0x+16000/1.0x^2+16000/1.0x^3+16000/1.0x^4

41,424=16000[1/1.0x+1/1.0x^2+1/1.0x^3+1/1.0x^4]

41,424/16000=[1/1.0x+1/1.0x^2+1/1.0x^3+1/1.0x^4]

[1/1.0x+1/1.0x^2+1/1.0x^3+1/1.0x^4]=2.589

a.Hence present value factor for annuity=2.589

b.Looking at present value annuity factor (20%,4 years);we get irr=20%